JKL Co. issues zero coupon bonds on the market at a price of $348 per bond. Each bond has a face value of $1,000 payable at maturity in 14 years. What is the yield to maturity for these bonds (in percent)? Answer to two decimals. Assume 1,000 par value and semi annual compounding
Answer #1

Current price=Par value/(1+Yield to maturity/2)^(2*time period)

348=1000/(1+Yield to maturity/2)^(2*14)

(1+Yield to maturity/2)^28=(1000/348)

1+Yield to maturity/2=(1000/348)^(1/28)

Yield to maturity=[1.03841791-1]*2

=7.68%(Approx).